# Learning not cheating: AI assistance can enhance rather than hinder skill development

### v2

@@ Discussion
Contrary to the expectations of lay forecasters (Study 1), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 3). Specifically, participants who had practiced with AI wrote (without AI assistance) cover letters that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later. Learning gains were not the result of greater effort; in fact, participants who’d practiced with AI expended less effort during practice than those who’d practiced alone. Instead, these gains can be explained by exposure to a high-quality, just-in-time personalized example: participants who merely viewed an AI-generated example cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 3).
While not obvious to forecasters, teachers (11), and the general population (8), the benefits of seeing AI-generated examples are consistent with prior research on how people learn. In addition to the literature on worked examples mentioned earlier, research has shown that humans are especially adept at observing, imitating, and learning from others (28, 29, 30). Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (31).
@@
Three promising directions for future research are worth highlighting. First, it remains to be seen whether learning from AI examples occurs in domains other than writing, currently the modal use case for AI tools in the workplace (25) . Writing may be particularly suited to learning by example because at a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting). In other domains, merely observing a solution may be less informative. For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it. Studies of AI tutors in these domains have, perhaps not surprisingly, found null or negative effects (23, 24).
Second, additional research is needed to explore moderators of learning from AI tools. Certain strategies for interacting with AI may bolster their effectiveness. For example, experimental research suggests that learners benefit more from AI explanations for math problems if they first try to solve them on their own (21). Similarly, correlational evidence that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22). On the other hand, other factors could minimize learning benefits. In our experimental paradigm, participants practiced for as long as they wanted, with the foreknowledge that their skills would subsequently be tested (and rewarded monetarily) without access to AI. During practice, therefore, participants were incentivized to prioritize gains in acquired skill over performance in the moment. In real-world settings, there is often time pressure and competing incentives around performance and learning, which we speculate would reduce the learning gains associated with practicing with AI.
Third, in our experimental paradigm, participants who interacted with the AI tool did so only once. It is common, however, to use AI tools repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Consider, for instance, the game of Go. The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (32). Future research, ideally in field settings, is needed to establish the long-term benefits and costs of relying on AI tools.
@@ Conclusion
Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development. Although it reduced the effort users invested in practicing, the AI writing tool nevertheless accelerated skill development. It accomplished this by providing high-quality, just-in-time, personalized examples of excellent writing. The underappreciated efficacy of timely and tailored examples has practical implications for the design of AI tools. Many AI tools designed to support learning are explicitly programmed not to “give away” answers. It may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
Decades before the advent of generative AI, the legendary UCLA baseball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (33). Few learners have access to the best human teachers, coaches, and mentors, but generative AI now makes it possible to learn from personalized, just-in-time demonstrations tailored to any domain. In doing so, AI has the potential not only to boost productivity but also to democratize opportunities to build human capital at scale.
@@ Methods
All our studies were pre-registered. See pre-registrations for Study 1 at https://aspredicted.org/x9mm-7qwp.pdf.
Study 2 was run twice because of a technical issue that caused imbalanced missing data issue the first time it was run. The findings in this sample were consistent with the ones reported here. Pre-registration is available at https://aspredicted.org/4sw4-mpny.pdf. The version of Study 2 presented above was pre-registered https://aspredicted.org/xyyn-gmzc.pdf.
Pre-registartion for Study 3 is available at https://aspredicted.org/3mty-fcfy.pdf. The one-day follow-up for Study 3 was collected in three batches. We pre-registered batch 2 https://aspredicted.org/5jcx-bhg9.pdf, but report pooled results in the main text. Details on the separate batches are available in Supplementary Information Section C5.
@@
Participants in Study 1 (N = 150) were predominantly female (n = 93, 62%), and ranged in age from 21 to 81 (M = 38.4, SD = 12.2). They were predominantly white (75%). A small proportion were students (13%), and most were employed (62%)
In Study 2, the sample was more evenly split between men (46%) and women (52%), and ranged in age from 18 to 82 (M = 36.0, SD = 12.5). Over half of the sample (58%) was white, with the rest being comprised by Black (33%), latino (6%), and asian (5%). 77% had college degrees. Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before).
Study 3 had similar prooportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6). Over half of the sample (64%) was white, with the rest being comprised of Black (24%), Latino (8%), and Asian (6%) participants. Most participants (74%) had college degrees. Most participants (91%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (40% had tried them, but hardly ever used them, 42% used them at least a few times per week, and 18% had never used AI assistants before).
@@
Participants first saw an introductory screen about what they were about to do. They then completed a brief questionnaire where they reported their demographics, their experience with AI writing tools, their motivation to improve their writing, and their perceived writing skill. They then completed a 2-minute pre-test in which they saw a poorly worded cover letter and were asked to improve it. After this, all participants completed a lesson about the five principles of effective writing. They were then randomized to the practice condition, or skipped ahead if assigned to the no-practice control. During the practice phase, participants rewrote a different cover letter, or (in the see example condition) generated an AI rewritten version of that letter. This example was not explicitly labeled as AI-generated. Then, participants saw the text they submitted (or the AI-generated example), and below it, AI-generated feedback highlighting one way in which this letter could be improved. See more information about the feedback procedure in Supplementary Information section A3. Immediately after seeignt this feedback screen, all participants then completed two questions, reporting how much they had learned and how hard they had worked on the task so far. They then edited a new cover letter in an incentivized, 7-minute test without the help of any AI tools. To minimize the possibility of cheating, we used custom JavaScript to restrict copy-pasting functionality. Finally, participants were invited to see optional feedback and were asked if they had used any outside resources during the test.
@@
As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (34, 35). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Table S6 and S16.
To validate these ratings, the first author and a trained research assistant took a sample of n = 30 cover letters from Study 2, and rated them on the 5 principles. The average of these two ratings correlated more highly with the computer ratings (r = .83, p < .001), than the average interrater reliability (r = .69, p < .001).
To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters. We find that GPT ratings correlate highly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6 and S16), suggesting that our effects are not explainable by same-model bias.
@@
As per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness). We used logistic regression to predict whether participants chose to see optional feedback for their test from condition, controlling for pretest score and baseline characteristics. Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1. When correcting for multiple comparisons in exploratory moderation analyses, we use the Benjamini-Hochberg correction (36).
C Results Study 3
@@
- C2 AI examples improve writing skill
- C3 Seeing AI examples was less effortful
- C4 Seeing an AI example did not discourage motivation for future learning
- C5 The effects of seeing an AI example persist
- C6 Seeing AI examples was equally effective across subgroups
@@ Appendix A Additional methods
After participants completed the procedure outlined in Figure 2, we had three writing samples for participants who practiced with or without AI (one for each phase of the experiment: pretest, practice, test), and two writing samples for participants who did not practice, or simply saw an AI generated example. We used GPT-4o to rate these texts for five writing principles. Each text and rating was completed independently of each other (i.e., the model had no memory of seeing that text before or of having rated it for any of the other writing principles). For robustness checks, we also used Anthropic’s Claude Haiku.
@@
Claude and GPT-4o ratings were positively correlated, with pretest ratings being less so. In Study 2, the correlations were .38, .78, .65, and .67 for pretest, practice, test, and followup, respectively. All p-values were below .001.
@@
Prolific participants were shown pairs of cover letters sampled from different conditions. They were asked to “Imagine you’re hiring a social media manager for your company; which cover letter would make you more likely to offer an interview to the candidate? Choose one.”. Each cover letter was compared to at least three other letters, sampled uniformly at random from the other two conditions. Most letters were compared against 3 or 4 other letters, min, max, med (plot?). For each cover letter we calculated the relative likelihood of it securing a hypothetical interview, defined as the total number of times that letter was preferred, divided by the total number of contests for that cover letter.
As shown in Figure S2, the relative likelihood of an interview correlated positively with the AI-generated measure of writing skill.
| (1) | (2) | (3) | (4) | |
| (Intercept) | *** | ** | ||
| No practice | *** | *** | ||
| Practice wo AI | ** | |||
| Practice w AI | ** | |||
| See AI example | *** | |||
| Precision () | *** | *** | ||
| Symmertry (Log()) | *** | |||
| Demographic and baseline performance covariates | No | Yes | No | Yes |
| Num.Obs. | ||||
| AIC | ||||
| BIC | ||||
| Log.Lik. | ||||
| RMSE |
@@
In Studies 2 and 3, participants received feedback for their submissions. The feedback was displayed immediately after the practice cover letter submission. The feedback page read: “Here is the email”, then reproduced the participants submission verbatim, then read “Here is one way in which it could be made better.” The feedback was personalized and created by GPT-4o model.
The feedack prompt is shown in Figure S3.
Appendix B Results Study 2
B1 Randomization, Balance, and Missingness

### v3

@@ Discussion
Contrary to lay forecasters’ expectations (Studies 1 and 3), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 5), more than those who had access to Google cover letter examples, and even more than those who received personalized feedback from professional human editors (Study 4). Specifically, participants who had practiced with AI, subsequently wrote cover letters (without AI assistance) that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later. Learning gains were not the result of greater effort; in fact, participants who had practiced with AI expended less effort during practice than those who had practiced alone. Instead, these gains were at least in part explained by exposure to high-quality, just-in-time personalized examples: participants who merely viewed an AI-generated cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 5).
Study 1 forecasters who believed AI would harm learning cited reduced effort as the mechanism—and prior research supports this concern (28, 29). All things equal, the more effort a learner invests in a given practice task, the more they learn. So why did practicing with AI improve writing skill more than practicing alone, even while reducing both the quantity and quality of effort? Likewise, why did writers who googled how to compose a cover letter work longer and harder than writers who practiced with AI, yet learn less? We propose a third mechanism that offsets decrements in human effort: the utility of the practice task itself. See Figure 8. Perhaps the excellent personalized examples provided by AI are analogous to demonstrations of an Olympic coach. The higher the utility of the practice task, the faster the learning.
While not immediately apparent to forecasters, teachers (11), and the general population (8), the pedagogical benefits of viewing AI-generated examples are consistent with prior research on learning. Research has shown that humans are especially adept at observing, imitating, and learning from others (30, 31, 32). Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (33).
@@
Three promising directions for future research are worth highlighting. First, it is not clear whether the observed benefits of AI examples would be as large in domains other than professional writing. At a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting). In other domains, however, merely observing a solution may be less informative. For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it.
Second, additional research is needed to explore moderators of learning from AI tools. Certain metacognitive strategies may enhance the benefits of interacting with AI. For example, experimental research suggests that learners benefit more from AI explanations for math problems if they first try to solve them on their own (21). Similarly, correlational evidence suggests that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22). Conversely, other factors may minimize learning benefits. In our experimental paradigm, participants practiced for as long as they wanted, with the foreknowledge that their skills would subsequently be tested (and rewarded monetarily) without access to AI. During practice, therefore, participants were incentivized to prioritize gains in acquired skill over performance in the moment. In real-world settings, there is often time pressure and competing incentives around performance and learning, which we speculate would reduce the learning gains associated with practicing with AI. More broadly, we believe concerns about AI as a crutch are justified when AI displaces effort without increasing task utility.
Third, in our experimental paradigm, participants interacted with the AI tool only once. It is common, however, to use AI tools repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Consider, for instance, the game of Go. The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (34). Future research, ideally in field settings, is needed to establish the long-term benefits and costs of using AI tools.
@@ Conclusion
Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development. Although the AI writing tool reduced the effort users invested in practicing, it nevertheless accelerated their skill development. The underappreciated efficacy of timely and tailored examples carries practical implications: Many AI tools designed to support learning are explicitly programmed not to “give away” answers. However, it may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
Decades before the advent of generative AI, the legendary UCLA basketball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (35). Few learners have access to the best human teachers, coaches, and mentors, but generative AI now makes it possible to learn from personalized, just-in-time demonstrations tailored to any domain. In doing so, AI has the potential not only to boost productivity but also to democratize opportunities for building human capital at scale.
@@ Methods
All studies were preregistered on https://aspredicted.org. Accordingly, the analyses presented in the main text were also pre-registered, with exceptions noted below.
Study 1 was preregistered (#191800). The moderation analyses were not pre-registered.
Study 2 was pre-registered (#201140). The pairwise comparisons analysis was pre-registered separately (#205316). After preregistration, we considered a keystrokes-per-minute metric of effort intensity rather than duration; thus, this analysis was not preregistered.
Study 3 was pre-registered (#236487). Analyses on predicted effectiveness of AI vs. human feedback and the mediation analysis were not preregistered.
Study 4 was pre-registered (#239157). The pairwise comparisons analysis was pre-registered separately (#243115). The keystrokes-per-minute metric of effort was not pre-registered.
Study 5 was pre-registered (#197704). The one-day follow-up was collected in three batches: a pilot, followed by second and third batches. We pre-registered batch 2 (#199451) but report pooled results in the main text, with separate batch details available in Supplementary Information Section F5. The pairwise comparisons analysis was pre-registered separately (#205315). The keystrokes-per-minute metric of effort was not pre-registered.
@@
Participants in Study 1 were predominantly female (n = 93, 62%), with ages ranging from 21 to 81 (M = 38.4, SD = 12.2). They were predominantly white (75%). A small proportion were students (13%), and most were employed (62%).
In Study 2, the sample was more evenly split between men (46%) and women (52%), ranging in age from 18 to 82 (M = 36.0, SD = 12.5). Over half of the sample (58%) was White, with the rest comprising Black (33%), Latino (6%), and Asian (5%) participants. Most participants (77%) had college degrees. Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before).
Participants in Study 3 were evenly split between men (52%) and women (48%), and their ages ranged from 18 to 75 (M = 42.4, SD = 14.4). They were predominantly white (67%). A small proportion were students (19%), and most were employed (56%).
In Study 4, the sample consisted of relatively fewer men (38%) than women (61%), and ranged in age from 19 to 92 (M = 41.5, SD = 13.8). Over half of the sample (74%) was White, with the rest comprising Black (16%), Latino (8%), and Asian (7%) participants. Most participants (74%) had college degrees. Most participants (94%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants: 17% had never used AI assistants, 40% had tried them but hardly ever used them, and 43% used them at least weekly.
Study 5 had similar proportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6). Over half of the sample (64%) was White, with the remainder comprising Black (24%), Latino (8%), and Asian (6%) participants. Most participants (74%) had college degrees. Most participants (91%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (40% had tried them, but hardly ever used them, 42% used them at least a few times per week, and 18% had never used AI assistants before).
@@
All studies followed the same general structure with study-specific variations. Participants first saw an introductory screen and completed a brief questionnaire reporting demographics, prior experience with AI writing tools, motivation to improve their writing, and perceived writing skill. They then completed a short pre-test (2 minutes) in which they edited a poorly written cover letter, followed by a brief lesson on the five principles of effective writing.
In Studies 2 and 5, participants were then randomized to a practice condition (or skipped practice in the no-practice control). During practice, participants rewrote a new cover letter or, in the example-only condition, viewed an AI-rewritten version of the letter. This example was not explicitly labeled as AI-generated. After practice, participants saw either their own rewritten letter or the AI-generated example, along with AI-generated feedback highlighting one suggested improvement (see Supplementary Information, Section A3). Immediately after this feedback, participants rated how much they had learned and how hard they had worked so far. They then completed a 7-minute incentivized test of writing skill by editing a new cover letter without access to AI tools. To minimize the possibility of cheating, we used custom JavaScript to restrict copy-pasting functionality. Finally, participants could request optional feedback and were asked whether they had used outside resources during the test. A small percentage of participants (2.9%) admitted to doing so; as per our preregistration, these participants are included in analyses, though results excluding them are consistent with our main interpretation (see Tables S6, S18, and S27).
In Study 4, we adapted the procedure to compare AI practice with two ecologically valid human alternatives: feedback from professional editors and guidance from Google Search. Participants completed the same baseline questionnaire, pretest, and writing lesson. They then edited a new cover letter without any assistance and were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional editors, (3) practice with a Google Search for cover letter examples and tips. In the second condition condition, participants’ cover letters were forwarded to editors recruited in partnership with The Journalists’ Resource at Harvard Kennedy School, who provided both written feedback and a rewritten version of the letter. The following day, participants were asked to revise their original cover letter using the resources from their assigned condition. All participants then completed a final writing test by rewriting a new cover letter without assistance, followed by an exit questionnaire.
@@
As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (36, 37). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Tables S6, S18, and S27.
To validate these ratings, the first author and a trained research assistant took a sample of n = 100 cover letters from Study 2, and rated them on the 5 principles. The average of these two ratings correlated about as highly with the computer ratings (r = .70, p < .001) as they did with the average interrater reliability (r = .74, p < .001). These ratings had high internal consistency (Cronbach’s s = 0.81, 0.78, and 0.83 for studies 2, 4, 5, respectively.)
To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters. We find that GPT ratings correlate strongly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6, S18, and S27), suggesting that our effects are not explainable by same-model bias.
Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. Cover letters more likely to secure an interview obtained higher AI ratings of writing skill (rs = .29, .35, and .28, ps < .001, for Study 2, 4 and 5, respectively).
@@
As per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness). We used logistic regression to predict whether participants chose to see optional test feedback from condition, controlling for pretest score and baseline characteristics. Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1. When correcting for multiple comparisons in exploratory moderation analyses, we used the Benjamini-Hochberg correction (38).
C Results Study 2
@@
- C2 AI practice improved writing skill
- C3 AI practice was less effortful
- C4 AI practice did not discourage motivation for future learning
- C5 The benefits of practicing with AI were just as large a day later
- C6 AI practice was equally effective across subgroups
- C7 Pairwise comparisons
- D Results Study 3
E Results Study 4
- E1 Randomization, Balance, and Missingness
- E2 Details on professional editors
- E3 Details on participants’ search behavior
- E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
- E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
- E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
- E7 Results were not moderated by individual differences
- E8 Pairwise comparisons
F Results Study 5
- F1 Randomization, Balance, and Missingness
- F2 AI examples improve writing skill
- F3 Seeing AI examples was less effortful
- F4 Seeing an AI example did not discourage motivation for future learning
- F5 The benefits of seeing an AI example were just as large a day later
- F6 Seeing AI examples was equally effective across subgroups
- F7 Pairwise Comparisons
@@ Appendix A Additional methods
After participants completed the procedure outlined in Figure 2, we had three writing samples for participants who practiced with or without AI (one for each phase of the experiment: pretest, practice, test), and two writing samples for participants who did not practice, or simply saw an AI-generated example. We used GPT-4o to rate these texts for five writing principles. Each text and rating was completed independently of each other (i.e., the model had no memory of seeing that text before or of having rated it for any of the other writing principles). For robustness checks, we also used Anthropic’s Claude Haiku.
@@
Claude and GPT-4o ratings were positively correlated, with pretest ratings being less so. In Study 2, the correlations were .38, .78, .65, and .67 for pretest, practice, test, and follow-up, respectively. All p-values were below .001.
@@
Prolific participants were shown pairs of cover letters sampled from different conditions. They were asked to “Imagine you’re hiring a social media manager for your company; which cover letter would make you more likely to offer an interview to the candidate? Choose one.” Each cover letter was compared to at least three other letters, sampled uniformly at random from the other two conditions. Most letters were compared against 3 or 4 other letters. For each cover letter, we calculated the relative likelihood of it securing a hypothetical interview, defined as the total number of times that letter was preferred, divided by the total number of contests for that cover letter.
@@
In Studies 2 and 3, participants received feedback for their submissions. The feedback was displayed immediately after the practice cover letter submission. The feedback page read: “Here is the email, then reproduced the participant’s submission verbatim, then read “Here is one way in which it could be made better.” The feedback was personalized and created by the GPT-4o model.
The feedback prompt is shown in Figure S2.
Appendix B Results Study 1
Table S2 reports results from a logistic regression predicting the probability of answering correctly as a function of demographic and background characteristics. Odds ratios (ORs) are displayed along with their standard errors, Wald z statistics, and p-values. The model includes age, sex, ethnicity, student status, employment, and prior AI experience as predictors.
| Predictor | OR | SE | z | p-value |
| Prior AI experience | 0.718 | 0.199 | -1.664 | 0.096 |
| Age (years) | 1.014 | 0.021 | 0.675 | 0.500 |
| Male (vs. Female) | 1.035 | 0.467 | 0.073 | 0.942 |
| White (vs. Other) | 0.966 | 0.509 | -0.068 | 0.946 |
| Student (vs. Not) | 1.120 | 0.618 | 0.183 | 0.855 |
| Employment status | ||||
| Other | 0.770 | 0.687 | -0.382 | 0.703 |
| Part-time | 0.663 | 0.601 | -0.684 | 0.494 |
| Unemployed (seeking) | 0.645 | 0.593 | -0.738 | 0.461 |
Appendix C Results Study 2
C1 Randomization, Balance, and Missingness